# Analysis of mean field games via Fokker-Planck-Kolmogorov equations:

### v1

@@
Abstract We study mean field games with unbounded coefficients. The existence of a solution is proved. We propose a new approach based on Fokker–Planck–Kolmogorov equations, the Ambrosio–Figalli–Trevisan superposition principle and a priory estimates with Lyapunov functions.
@@ 1. Introduction
The main result presented in Theorem 2.1 is the existence of a solution to the problem (P). This problem arises in stochastic mean field games which have the following structure. Let us fix a measure $\sigma=\sigma_{t}\,dt$ such that the mapping $t\mapsto\sigma_{t}$ with values in the space of probability measure on $\mathbb{R}^{d}$ is continuous with respect to the weak topology. Solve the optimal control problem
@@
Denote by $\Phi(\sigma)$ the set of all measures $\eta=\eta_{t}\,dt$ such that there exists a solution $(X_{t},U_{t})$ of this optimal control problem and $\eta_{t}={\rm Law}\bigl{(}X_{t}\bigr{)}$ for all $t\in[0,T]$. A mean field game solution $\mu$ is a fixed point of the mapping $\sigma\mapsto\Phi(\sigma)$ that is $\mu\in\Phi(\mu)$. Mean field games describe the Nash equilibria in games with an infinite number of agents. In this interpretation the solution $\mu$ represents the distribution of an infinity of agents state processes and the corresponding $X_{t}$ describes the state process dynamics of a single representative agent. Mean field games are used as approximations to the Nash equilibria in stochastic games with many players. The solution $(u,\mu)$ of the problem (P) and the Ambrosio–Figally–Trevisan superposition principle together with the convexity assumptions allow us to solve the mean field game. This is discussed in Corollary 2.1. Note that if $\mu_{t}={\rm Law}\bigl{(}X_{t}\bigr{)}$ and $U_{t}=u(X_{t},t)$ for some Borel function $u$, then
@@
Note that if the coefficients $A$, $b$, $Q$ and the functions $f$, $g$ do not depend on $\mu$, then the problem (P) is a control problem for linear Fokker–Planck–Kolmogorov equations (see, for instance [2], [3], [4]). If $f=g=0$ and $Q=0$, then the problem (P) is the Cauchy problem for the nonlinear Fokker–Planck–Kolmogorov equation (see, for instance [10], [16], [30], [35]). We stress that the problem (P) is not a control problem for nonlinear Fokker–Planck–Kolmogorov equations (see [26]). The case $A=0$ corresponds to the deterministic mean field games which are discussed in papers [5], [20], [22]. Finally, note that mean field games for more general Markov process are considered in works [6], [27], [36].
@@ 2. Main results
Let $T>0$, $V\in C^{2}(\mathbb{R}^{d})$, $V\geq 0$ and $\lim_{|x|\to+\infty}V(x)=+\infty$. For example one can take the function $V(x)=(1+|x|^{2})^{p/2}$, where $p>0$. The bounded Borel measure $\mu$ on $\mathbb{R}^{d}$ is called a probability measure if $\mu\geq 0$ and $\mu(\mathbb{R}^{d})=1$. We say that the bounded Borel measure $\mu$ on $\mathbb{R}^{d}\times[0,T]$ is given by a family of Borel measures $(\mu_{t})_{t\in[0,T]}$ if for every Borel set $E$ the mapping $t\mapsto\mu_{t}(E)$ is Borel measurable and for every bounded Borel function $\eta$ the equality
@@
We shall say that measures $\mu^{n}=\mu^{n}_{t}\,dt$ in the set $\mathcal{M}(V)$ converge $V$ – weakly to a measure $\mu=\mu_{t}\,dt$ in $\mathcal{M}(V)$ if for all $t\in[0,T]$ one has the equality
@@
for every continuous function $\zeta$ on $\mathbb{R}^{d}$ with $\lim_{|x|\to+\infty}\zeta(x)/V(x)=0$.
@@
Below we often use the inequality $pv\leq h(p)+h(v)$.
@@
on $U\times\mathbb{R}^{d}\times[0,T]$ and on $\mathbb{R}^{d}$ respectively.
@@
$\bf(H3.3)$ For all open balls $B\subset\mathbb{R}^{d}$, $B_{1}\subset\mathbb{R}^{d_{1}}$ and every number $R>0$ the VV–weak convergence of measures $\mu^{n}\in\mathcal{M}_{R,M}(V)$ to a measure $\mu\in\mathcal{M}_{R,M}(V)$ implies the equality
@@
Let $V(x)=1+|x|^{m}$, $W(x)=|x|^{p}$ and $h(v)=Cv^{2}$, where $m\geq 2$, $1\leq p<m$ and $C>0$. Suppose that there exist positive numbers $C_{1}$, $C_{2}$, $C_{3}$, $C_{4}$, $C_{5}$, $C_{6}$ and $\varepsilon$ such that for every $x,y\in\mathbb{R}^{d}$, all $t\in[0,T]$ and every $\mu\in\mathcal{M}(V)$ we have
@@
$\|\sqrt{A(x,t,\mu)}\|\leq C_{1}\bigl{(}1+|x|\bigr{)},\quad\|Q(x,t,\mu)\|\leq C_{1}\bigl{(}1+|x|^{1-\varepsilon}\bigr{)},$
@@
For instance, this condition holds for $b(x,t,\mu)=-x|x|^{m-1}+b_{0}(t,\mu)$, where $b_{0}$ is bounded vector fields.
@@
$h^{*}(|Q^{\top}(x,t,\mu)\nabla V(x)|)=\frac{m^{2}}{4C}\|Q(x,t,\mu)\|^{2}|x|^{2m-2}\leq\frac{m^{2}C_{1}^{2}}{4C}|x|^{2m-2}+\frac{m^{2}C_{1}^{2}}{4C}|x|^{2m-1-\varepsilon}.$
Since $2m-1-\varepsilon<2m-1$, we arrive at the estimate
@@
$h^{*}(\|Q(x,t,\mu)\|)\leq\frac{C_{1}^{2}}{4C}\bigl{(}1+|x|^{1-\varepsilon}\bigr{)}^{2}\leq N_{3}V(x),$
@@
In general in order to construct functions $h$, $V$ and $W$ the following approach is suggested. Firstly, we find functions $h$ and $W$ such that the conditions (H3.1) and (H3.2) are fulfilled. Secondary, using the functions $h$ and $W$, we obtain a function $V$ such that $V\in C^{2}(\mathbb{R}^{d})$, $W\leq V$, $\lim_{|x|\to\infty}V(x)=+\infty$, $\lim_{|x|\to\infty}W(x)/V(x)=0$ and the condition (H2.1) is fulfilled. Thirdly, we find a function $\Theta$ such that the condition (H2.2) holds. Finally, we verify the remaining growth conditions.
@@
over all probability measures $\pi$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with projections $\mu$ and $\sigma$ on the factors. We consider the space $\mathcal{P}_{p}(\mathbb{R}^{d})$ with the distance $W_{p}$.
@@
Let $\overline{B}$ denote the closure of a ball $B$. Note that $\mathcal{M}_{R}(V)$ is a compact set in $\mathcal{P}_{p}(\mathbb{R}^{d})$. It follows that $\overline{B}\times\mathcal{M}_{R}(V)$ is a compact set and for every $t\in[0,T]$ the mapping $(x,\eta)\mapsto\Psi(x,t,\eta)$ is uniformly continuous on $\overline{B}\times\mathcal{M}_{R}(V)$. Finally, note that the VV–weak convergence of measures $\mu^{n}=\mu^{n}_{t}\,dt\in\mathcal{M}_{R}(V)$ to a measure $\mu=\mu_{t}\,dt\in\mathcal{M}_{R}(V)$ implies for every $t\in[0,T]$ the equality $\lim_{n\to\infty}W_{p}(\mu_{t}^{n},\mu_{t})=0$. ∎
@@
Let $\overline{B}$ denote the closure of a ball $B$. Since $\overline{B}\times[-C_{\zeta}R,C_{\zeta}R]$ is a compact set the function $\Phi(x,t,r)$ is uniformly continuous in $(x,r)$ on $\overline{B}\times[-C_{\zeta}R,C_{\zeta}R]$. Therefore the equality
@@ 3. Auxiliary results
Similarly one can define the weak topology on the space of bounded Borel measures on $\mathbb{R}^{d}\times[0,T]$ and this topology is metrizable on the set of nonnegative measures $\mu$ satisfying the equality $\mu(\mathbb{R}^{d}\times[0,T])=T$. The sequence of probability measures $\mu^{n}$ on $\mathbb{R}^{d}$ is tight if for every $\varepsilon>0$ there exists a compact set $K_{\varepsilon}\subset\mathbb{R}^{d}$ such that $\mu^{n}(K_{\varepsilon})\geq 1-\varepsilon$ for all $n$. Let $\mathcal{V}$ be a continuous nonnegative function with $\lim_{|x|\to\infty}\mathcal{V}(x)=+\infty$ and $\mu^{n}$ be a sequence of probability measures. The estimate
$\sup_{n}\int_{\mathbb{R}^{d}}\mathcal{V}(x)\mu_{n}(dx)<\infty$
@@
Let $\mu^{n}$ be a sequence of probability measures on $\mathbb{R}^{d}$. If for every function $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ and every number $\varepsilon>0$ there exists a number $N$ such that the inequality
@@
holds for all $n,k>N$, then the measures $\mu^{n}$ weakly converge to some probability measure $\mu$. It is easily seen that the above condition implies the tightness of the sequence $\mu^{n}$. Thus the assertion follows from the Prokhorov theorem. Similarly probability measures $\mu^{n}$ on $\mathbb{R}^{d}$ converge weakly to a probability measure $\mu$ if the equality
@@
The weak topology on the space of measures is discussed in [12] and [11, Chapter 8].
@@
We need a modification of the estimates from [14, Theorem 7.1.1] and [15, Lemma 2.2]. Assume that there exist a nonnegative function $\mathcal{V}\in C^{2}(\mathbb{R}^{d})$ and nonnegative Borel function $\mathcal{W}$ on $\mathbb{R}^{d}\times[0,T]$ such that
$\lim_{|x|\to\infty}\mathcal{V}(x)=+\infty,\,\mathcal{L}\mathcal{V}(x,t)\leq\mathcal{W}(x,t)+C\mathcal{V}(x),\,\int_{0}^{T}\int_{\mathbb{R}^{d}}\mathcal{W}(x,t)\mu_{t}(dx)\,dt<\infty,$
@@
then $\mu_{t}(\mathbb{R}^{d})=1$ and $\|\mathcal{V}\|_{L^{1}(\mu_{t})}\leq C^{\prime}$ for almost all $t\in[0,T]$, where the constant $C^{\prime}$ does not depend on $t$. According to the [17, Proposition 4.1], one can redefine the solution $\mu_{t}$ in such a way that the equality (*) is fulfilled for all $t\in[0,T]$. Note that for every $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ there exists a number $C(\psi)>0$ such that the estimate
@@
Using the inequalities from the conditions (H3.1), (H3.2) and taking into account Remark 3.2, we obtain the estimate
@@
$L_{\sigma,u}V(x,t)\leq h(|u(x,t)|)+C_{L}V(x)+C_{L}\int_{\mathbb{R}^{d}}V(y)\sigma_{t}(dy)+C_{L}\sup_{t\in[0,T]}\int_{\mathbb{R}^{d}}W(y)\sigma_{t}(dy)\leq\\ h(|u(x,t)|)+C_{L}V(x)+C_{L}\int_{\mathbb{R}^{d}}V(y)\sigma_{t}(dy)+C_{L}\beta_{V,W}(Re^{MT})Re^{MT}$
@@
$\varepsilon(1+C_{f}T+C_{g}\bigr{)}+\varepsilon C_{f}\frac{1}{M}e^{MT}+\varepsilon C_{g}e^{MT}\leq\frac{\gamma}{2}.$
@@
Hence the sequence
$\int_{\mathbb{R}^{d}}\psi\,d\mu^{n_{j}}_{t}$
converges for every function $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$. According to Remark 3.1, the measures $\mu^{n_{j}}_{t}$ converge weakly to a probability measure $\mu_{t}$. Since for every $j$ the integral of $V$ with respect to the measure $\mu^{n_{j}}_{t}$ is majorized by $Re^{Mt}$, the same estimate holds for the integral of $V$ with respect to the measure $\mu_{t}$. Thus we obtain the family of probability measures $(\mu_{t})_{t\in[0,T]}$. Note that for every $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ the estimate
@@
Let us prove (ii). Suppose that measures $\mu^{n}=\mu^{n}_{t}\,dt\in\mathcal{M}_{R,M}(V)$ converge weakly to a measure $\mu=\mu_{t}\,dt\in\mathcal{M}_{R,M}(V)$. It suffices to show that for every $t\in[0,T]$ the sequence $\mu^{n}_{t}$ converges weakly to the measure $\mu_{t}$. Note that every sequence of natural numbers $n_{k}$ posses a subsequence $n_{k_{j}}$ such that for every $t$ the measures $\mu^{n_{k_{j}}}_{t}$ converge weakly to a probability measure $\widetilde{\mu}_{t}$. In addition the mapping $t\mapsto\widetilde{\mu}_{t}$ is continuous. Since $\mu_{t}\,dt=\widetilde{\mu}_{t}\,dt$, for every bounded and continuous function $\varphi$ the equality
@@
This statement is known but for the reader’s convenience we give a brief proof. We may assume that $|\eta|\leq 1$. Let $\varepsilon>0$. Since the projections of measures $P_{n}$ on $X$ converge weakly to the projection of the measure $P$ on $X$, there exists a compact set $K\subset X$ such that the estimate $P_{n}(K\times Y)\geq 1-\varepsilon$ holds for all $n$. According to the Scorza Dragoni theorem (see, for instance [37, Theorem 2]), there exists a compact set $C\subset Y$ such that $\pi(C)\geq 1-\varepsilon$ and the restriction of the function $\eta$ on $X\times C$ is continuous. Let $\widetilde{\eta}$ be a continuous function on $X\times Y$ such that $\widetilde{\eta}=\eta$ on $K\times C$ and $|\widetilde{\eta}|\leq 1$. We have
@@
Suppose that we are given complete separable metric spaces $X$ and $Y$. Let $\mu$ be a bounded Borel nonnegative measure on $X\times Y$. Denote by $\mu_{Y}$ the projection of $\mu$ on $Y$ that is $\mu_{Y}(E)=\mu(X\times E)$ for every Borel set $E$. By [11, Theorem 10.4.10] there exists a family of Borel probability measures $\mu^{y}$ on $X$ such that for every Borel set $E$ the mapping $y\mapsto\mu^{y}(E)$ is Borel measurable and for every Borel function $f\in L^{1}(\mu)$ the equality
@@
Let us consider a Borel measurable family $(\sigma^{y})_{y\in Y}$ on $X$ and a nonnegative bounded Borel measure $\eta$ on $Y$. According to [11, Theorem 10.7.2], we can define the bounded nonnegative Borel measure $\mu$ by the equality $\mu(dxdy)=\sigma^{y}(dx)\eta(dy)$ which means that for every Borel set $C$

### v2

# arXiv:2508.15029v2[math.AP] 27 Feb 2026

@@
Abstract We study mean field games with unbounded coefficients. The existence of a solution is proved. We propose a new approach based on Fokker–Planck–Kolmogorov equations, the Ambrosio–Figalli–Trevisan superposition principle, the method of doubling variables and a priory estimates with Lyapunov functions.
@@ 1. Introduction
The main result to be presented in Theorem 2.1 is the existence of a solution to the problem (P). This problem arises in stochastic mean field games which have the following structure. Let us fix a measure $\sigma=\sigma_{t}\,dt$ such that the mapping $t\mapsto\sigma_{t}$ with values in the space of probability measures on $\mathbb{R}^{d}$ is continuous with respect to the weak topology. Solve the optimal control problem
@@
Denote by $\Phi(\sigma)$ the set of all measures $\eta=\eta_{t}\,dt$ such that there exists a solution $(X_{t},U_{t})$ of this optimal control problem and $\eta_{t}={\rm Law}\bigl(X_{t}\bigr)$ for all $t\in[0,T]$. A mean field game solution $\mu$ is a fixed point of the mapping $\sigma\mapsto\Phi(\sigma)$, that is $\mu\in\Phi(\mu)$. Mean field games describe the Nash equilibria in games with an infinite number of agents. In this interpretation the solution $\mu$ represents the distribution of state processes an infinity of agents and the corresponding $X_{t}$ describes the state process dynamics of a single representative agent. Mean field games are used as approximations to the Nash equilibria in stochastic games with many players. The solution $(u,\mu)$ of the problem (P) and the Ambrosio–Figally–Trevisan superposition principle together with the convexity assumptions allow us to solve the mean field game. This will be discussed in Corollary 2.1. Note that if $\mu_{t}={\rm Law}\bigl(X_{t}\bigr)$ and $U_{t}=u(X_{t},t)$ for some Borel function $u$, then
@@
There are several ways to construct solutions to mean field games. The first approach is to solve the forward-backward problem which has the form of a system of Hamilton–Jacobi–Bellman and Fokker–Planck–Kolmogorov equations (see, for instance [7], [43]). In this case coefficients have a special structure, for instance, the diffusion matrix $A$ is non-degenerate and even constant and $b=-H_{p}(x,t,p)$, in particular $A$ and $b$ do not depend on $\mu$. Moreover, the function $f$ has the form $f(u,x,t,\mu)=l(u,x,t)+F(x,t,\mu)$. However, this approach allows to take into account the nonlinearity of local type in $\mu$ and non-smooth coefficients. The second approach is based on a probabilistic analysis of mean field games (see [24], [26], [40]). The probabilistic approach is developed in three directions: 1) the stochastic maximum principle (see [11], [24]), 2) the convergence of the Nash equilibria in a symmetric NN–player game to the mean field game limit (see [22], [31], [41]), 3) relaxed controls (see [26], [40]). The stochastic maximum principle and the convergence of Nash equilibria require restrictive smoothness conditions on the coefficients. More general results use relaxed controls. The idea is as follows: first, we construct a measure on the control set $U$ and then, using the convexity of the data and conditional measures, we obtain a control function $(x,t)\mapsto u(x,t)$. This idea is well–known in optimal control problems (see [30]).
One of the most general existence results is presented in [40], where there are two types of assumptions: (A) the coefficients are Lipschitzian and (C) the coefficients are bounded and continuous, the matrix $A$ is non-degenerate and the control set $U$ is compact. Note that the assumptions of the paper [40] only allow linear growth of the coefficients. In addition, if $U=\mathbb{R}^{d_{1}}$, then the assumption (A.2) from [40] implies that $Q(x,t,\mu)$ does not depend on $x$. We improve the results of [40] and allow nonlinear growth of the coefficients and more general dependence on $\mu$ (see examples 2.12.4). In [40] the relaxed control is a random process with values in the space of probability measures, the theory of martingale measures is applied. This approach is generalised with similar assumptions to mean field games with singular controls [33] and to mean field games with absorption [20]. We propose a different approach based on Fokker–Planck–Kolmogorov equations and not involving stochastic control theory. Moreover, a priori estimates with Lyapunov functions allow us to consider coefficients with arbitrary growth. Another improvement is that the coefficients $A$, $b$, $Q$ and the functions $f$, $g$ can depend on the entire measure $\mu=\mu_{t}\,dt$ rather than on the measure $\mu_{t}$ at time $t$. The connection between probabilistic solutions of Fokker–Planck–Kolmogorov equations and solutions to stochastic differential equations is based on the Ambrosio–Figalli–Trevisan superposition principle, which is now known under very general assumptions (see [16], [44]). Moreover, we use the method of doubling variables (see [9], [28]). References to the latest results concerning the theory of Fokker–Planck–Kolmogorov equations can be found in [15].
Note that if the coefficients $A$, $b$, $Q$ and the functions $f$, $g$ do not depend on $\mu$, then the problem (P) is a control problem for linear Fokker–Planck–Kolmogorov equations (see, for instance [2], [3], [4]). If $f=g=0$ and $Q=0$, then the problem (P) is a Cauchy problem for the nonlinear Fokker–Planck–Kolmogorov equation (see, for instance [8], [17], [32], [37]). We stress that the problem (P) is not a control problem for nonlinear Fokker–Planck–Kolmogorov equations (see [27]). The case $A=0$ corresponds to the deterministic mean field games which are discussed in papers [5], [21], [23]. Finally, note that mean field games for more general Markov processes are considered in works [6], [29], [38].
@@ 2. Main results
Let $T>0$, $V\in C^{2}(\mathbb{R}^{d})$, $V\geq 0$ and $\lim_{|x|\to+\infty}V(x)=+\infty$. For example, one can take the function $V(x)=(1+|x|^{2})^{p/2}$, where $p>0$. The bounded Borel measure $\mu$ on $\mathbb{R}^{d}$ is called a probability measure if $\mu\geq 0$ and $\mu(\mathbb{R}^{d})=1$. We say that the bounded Borel measure $\mu$ on $\mathbb{R}^{d}\times[0,T]$ is given by a family of Borel measures $(\mu_{t})_{t\in[0,T]}$ if for every Borel set $E$ the mapping $t\mapsto\mu_{t}(E)$ is Borel measurable and for every bounded Borel function $\eta$ the equality
@@
We will say that measures $\mu^{n}=\mu^{n}_{t}\,dt$ in the set $\mathcal{M}(V)$ converge $V$ – weakly to a measure $\mu=\mu_{t}\,dt$ in $\mathcal{M}(V)$ if for all $t\in[0,T]$ one has the equality
@@
for every continuous function $\zeta$ on $\mathbb{R}^{d}$ with $\lim_{|x|\to+\infty}\zeta(x)/V(x)=0$ (see also Remark 3.3).
@@
Below we often use the inequality $pv\leq h(p)+h^{*}(v)$. Note also that the function $h^{*}$ plays a significant role in the growth conditions on our coefficients $A$, $b$ and $Q$ while the function $h$ limits the growth of the function $f$ with respect to the control $u$.
@@
on $U\times\mathbb{R}^{d}\times[0,T]$ and on $\mathbb{R}^{d}$, respectively.
@@
$\bf(H3.3)$ For all open balls $B\subset\mathbb{R}^{d}$, $B_{1}\subset\mathbb{R}^{d_{1}}$ and every number $R>0$ the VV–weak convergence of measures $\mu^{n}\in\mathcal{M}_{R}(V)$ to a measure $\mu\in\mathcal{M}_{R}(V)$ implies the equality
@@
This stochastic differential equation is commonly referred to as the granular media equation or the McKean–Vlasov equation. In the last two decades this equation and the corresponding nonlinear Fokker–Planck–Kolmogorov equation have been being studied intensively. The results on the existence and uniqueness of the solution can be found in [17], [49], [50]. We stress that nonlinear Fokker–Planck–Kolmogorov equations with coefficients rapidly increasing at infinity appear in many physical models with unbounded potentials (see [32]). Moreover, in socio-economics models and in traffic models linear and nonlinear Fokker–Planck–Kolmogorov equations are often considered on a half-line or on an interval with diffusion coefficients degenerating at the boundary. By a change of variables such equations are transformed to equations on the line with coefficients rapidly increasing at infinity. Theorem 2.1 and Corollary 2.1 allow us to consider mean field games with coefficients rapidly increasing at infinity.
@@
Let $V(x)=1+|x|^{m}$, $W(x)=|x|^{p}$ and $h(v)=Cv^{2}$, where $m\geq 2$, $1\leq p<m$ and $C>0$. Suppose that there exist number $\varepsilon\in(0,1/2)$ and positive numbers $C_{1}$, $C_{2}$, $C_{3}$, $C_{4}$, $C_{5}$, $C_{6}$ such that for every $x,y\in\mathbb{R}^{d}$, all $t\in[0,T]$ and every $\mu\in\mathcal{M}(V)$ we have
@@
$\|\sqrt{A(x,t,\mu)}\|\leq C_{1}\bigl(1+|x|\bigr),\quad\|Q(x,t,\mu)\|\leq C_{1}\bigl(1+|x|^{\varepsilon}\bigr),$
@@
For instance, this condition holds for $b(x,t,\mu)=-x|x|^{m-1}+b_{0}(t,\mu)$, where $b_{0}$ is a bounded vector field.
@@
$h^{*}(|Q^{\top}(x,t,\mu)\nabla V(x)|)=\frac{m^{2}}{4C}\|Q(x,t,\mu)\|^{2}|x|^{2m-2}\leq\frac{m^{2}C_{1}^{2}}{2C}|x|^{2m-2}+\frac{m^{2}C_{1}^{2}}{2C}|x|^{2m-2+2\varepsilon}.$
Since $2m-2+2\varepsilon<2m-1$, we arrive at the estimate
@@
$h^{*}(\|Q(x,t,\mu)\|)\leq\frac{C_{1}^{2}}{4C}\bigl(1+|x|^{\varepsilon}\bigr)^{2}\leq N_{3}V(x),$
@@
In general, in order to construct functions $h$, $V$ and $W$ the following approach is suggested. Firstly, we find functions $h$ and $W$ such that the conditions (H3.1) and (H3.2) are fulfilled. Secondary, using the functions $h$ and $W$, we obtain a function $V$ such that $V\in C^{2}(\mathbb{R}^{d})$, $W\leq V$, $\lim_{|x|\to\infty}V(x)=+\infty$, $\lim_{|x|\to\infty}W(x)/V(x)=0$ and the condition (H2.1) is fulfilled. Thirdly, we find a function $\Theta$ such that the condition (H2.2) holds. Finally, we verify the remaining growth conditions.
@@
over all probability measures $\pi$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with projections $\mu$ and $\sigma$ onto the factors (see [14]). We consider the space $\mathcal{P}_{p}(\mathbb{R}^{d})$ with the distance $W_{p}$.
@@
Let $\overline{B}$ denote the closure of a ball $B$. Note that $\mathcal{M}_{R}(V)$ is a compact set in $\mathcal{P}_{p}(\mathbb{R}^{d})$. It follows that $\overline{B}\times\mathcal{M}_{R}(V)$ is a compact set and for every $t\in[0,T]$ the mapping $(x,\eta)\mapsto\Psi(x,t,\eta)$ is uniformly continuous on $\overline{B}\times\mathcal{M}_{R}(V)$. Finally, note that the VV–weak convergence of measures $\mu^{n}=\mu^{n}_{t}\,dt\in\mathcal{M}_{R}(V)$ to a measure $\mu=\mu_{t}\,dt\in\mathcal{M}_{R}(V)$ implies the equality $\lim_{n\to\infty}W_{p}(\mu_{t}^{n},\mu_{t})=0$ for every $t\in[0,T]$. ∎
@@
Let $\overline{B}$ denote the closure of a ball $B$. Since $\overline{B}\times[-C_{\zeta}R,C_{\zeta}R]$ is a compact set, the function $\Phi(x,t,r)$ is uniformly continuous in $(x,r)$ on $\overline{B}\times[-C_{\zeta}R,C_{\zeta}R]$. Therefore the equality
@@ 3. Auxiliary results
Similarly one can define the weak topology on the space of bounded Borel measures on the space $\mathbb{R}^{d}\times[0,T]$ and this topology is metrizable on the set of nonnegative measures $\mu$ satisfying the equality $\mu(\mathbb{R}^{d}\times[0,T])=T$. The sequence of probability measures $\mu^{n}$ on $\mathbb{R}^{d}$ is tight if for every number $\varepsilon>0$ there exists a compact set $K_{\varepsilon}\subset\mathbb{R}^{d}$ such that $\mu^{n}(K_{\varepsilon})\geq 1-\varepsilon$ for all $n$. Let $\mathcal{V}$ be a continuous nonnegative function with $\lim_{|x|\to\infty}\mathcal{V}(x)=+\infty$ and $\mu^{n}$ be a sequence of probability measures. The estimate
$\sup_{n}\int_{\mathbb{R}^{d}}\mathcal{V}(x)\mu^{n}(dx)<\infty$
@@
Let $\mu^{n}$ be a sequence of probability measures on $\mathbb{R}^{d}$. Assume that $\mu^{n}$ is tight. If for every function $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ and every number $\varepsilon>0$ there exists a number $N$ such that the inequality
@@
holds for all $n,k>N$, then the measures $\mu^{n}$ weakly converge to some probability measure $\mu$. Assertion follows from the Prokhorov theorem and the fact that for every $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ the sequence $\int_{\mathbb{R}^{d}}\psi(x)\mu^{n}(dx)$ converges. Moreover, probability measures $\mu^{n}$ on $\mathbb{R}^{d}$ converge weakly to a probability measure $\mu$ whenever the equality
@@
The weak topology on the space of measures is discussed in [13] and [12, Chapter 8].
@@
We need a modification of the estimates from [15, Theorem 7.1.1] and [16, Lemma 2.2]. Assume that there exist a nonnegative function $\mathcal{V}\in C^{2}(\mathbb{R}^{d})$ and nonnegative Borel function $\mathcal{W}$ on $\mathbb{R}^{d}\times[0,T]$ such that
$\lim_{|x|\to\infty}\mathcal{V}(x)=+\infty,\quad\mathcal{L}\mathcal{V}(x,t)\leq\mathcal{W}(x,t)+C\mathcal{V}(x),\quad\int_{0}^{T}\int_{\mathbb{R}^{d}}\mathcal{W}(x,t)\mu_{t}(dx)\,dt<\infty,$
@@
then $\mu_{t}(\mathbb{R}^{d})=1$ and $\|\mathcal{V}\|_{L^{1}(\mu_{t})}\leq C^{\prime}$ for almost all $t\in[0,T]$, where the constant $C^{\prime}$ does not depend on $t$. According to the [18, Proposition 4.1], one can redefine the solution $\mu_{t}$ in such a way that the equality (*) is fulfilled for all $t\in[0,T]$. Note that for every $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ there exists a number $C(\psi)>0$ such that the estimate
@@
For every $\sigma\in\mathcal{M}_{R,M}(V)$ the inequality
$\int_{\mathbb{R}^{d}}V(x)\sigma_{t}(dx)\leq Re^{MT}$
holds for every $t\in[0,T]$. Taking into account the definition of the function $\beta_{V,W}$ from Remark 3.2 we obtain
$\sup_{t\in[0,T]}\int_{\mathbb{R}^{d}}W(x)\sigma_{t}(dx)\leq\beta_{V,M}(Re^{MT})Re^{MT}.$
Applying the inequalities from the conditions (H3.1) and (H3.2), we arrive at the estimate
@@
$L_{\sigma,u(x,t)}V(x,t)\leq h(|u(x,t)|)+C_{L}V(x)+C_{L}\int_{\mathbb{R}^{d}}V(y)\sigma_{t}(dy)+C_{L}\sup_{t\in[0,T]}\int_{\mathbb{R}^{d}}W(y)\sigma_{t}(dy)\leq\\ h(|u(x,t)|)+C_{L}V(x)+C_{L}\int_{\mathbb{R}^{d}}V(y)\sigma_{t}(dy)+C_{L}\beta_{V,W}(Re^{MT})Re^{MT}$
@@
$\varepsilon(1+C_{f}T+C_{g}\bigr)+\varepsilon C_{f}M^{-1}e^{MT}+\varepsilon C_{g}e^{MT}\leq\frac{\gamma}{2}.$
@@
According to Remark 3.1, the measures $\mu^{n_{j}}_{t}$ converge weakly to a probability measure $\mu_{t}$. Since for every $j$ the integral of $V$ with respect to the measure $\mu^{n_{j}}_{t}$ is majorized by $Re^{Mt}$, the same estimate holds for the integral of $V$ with respect to the measure $\mu_{t}$. Thus we obtain the family of probability measures $(\mu_{t})_{t\in[0,T]}$. Note that for every $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ the estimate
@@
Let us prove (ii). Suppose that measures $\mu^{n}=\mu^{n}_{t}\,dt\in\mathcal{M}_{R,M}(V)$ converge weakly to a measure $\mu=\mu_{t}\,dt\in\mathcal{M}_{R,M}(V)$. It suffices to show that for every $t\in[0,T]$ the sequence $\mu^{n}_{t}$ converges weakly to the measure $\mu_{t}$. Note that every sequence of natural numbers $n_{k}$ posses a subsequence $n_{k_{j}}$ such that for every $t$ the measures $\mu^{n_{k_{j}}}_{t}$ converge weakly to a probability measure $\widetilde{\mu}_{t}$. In addition, the mapping $t\mapsto\widetilde{\mu}_{t}$ is continuous. Since $\mu_{t}\,dt=\widetilde{\mu}_{t}\,dt$, for every bounded and continuous function $\varphi$ the equality
@@
This statement is known but for the reader’s convenience we give a brief proof. We may assume that $|\eta|\leq 1$. Let $\varepsilon>0$. Since the projections of measures $P_{n}$ on $X$ converge weakly to the projection of the measure $P$ on $X$, there exists a compact set $K\subset X$ such that the estimate $P_{n}(K\times Y)\geq 1-\varepsilon$ holds for all $n$. According to the Scorza Dragoni theorem (see, for instance, [39, Theorem 2] or [12, Theorem 7.14.26]), there exists a compact set $C\subset Y$ such that $\pi(C)\geq 1-\varepsilon$ and the restriction of the function $\eta$ on $X\times C$ is continuous. Let $\widetilde{\eta}$ be a continuous function on $X\times Y$ such that $\widetilde{\eta}=\eta$ on $K\times C$ and $|\widetilde{\eta}|\leq 1$. We have
@@
Suppose that we are given complete separable metric spaces $X$ and $Y$. Let $\mu$ be a bounded Borel nonnegative measure on $X\times Y$. Denote by $\mu_{Y}$ the projection of $\mu$ on $Y$ that is $\mu_{Y}(E)=\mu(X\times E)$ for every Borel set $E$. By [12, Theorem 10.4.10] there exists a family of Borel probability measures $\mu^{y}$ on $X$ such that for every Borel set $E$ the mapping $y\mapsto\mu^{y}(E)$ is Borel measurable and for every Borel function $f\in L^{1}(\mu)$ the equality
@@
Let us consider a Borel measurable family $(\sigma^{y})_{y\in Y}$ on $X$ and a nonnegative bounded Borel measure $\eta$ on $Y$. According to [12, Theorem 10.7.2], we can define the bounded nonnegative Borel measure $\mu$ by the equality $\mu(dxdy)=\sigma^{y}(dx)\eta(dy)$ which means that for every Borel set $C$